QUESTION 13 6.25 points S Assume that you hold a two stock portfolio. Use the following table to calculate the Price 24 annualized HPY of your portfolic. Initial Price Stock Weight 1/1/20131/1/2017 75 25 18 110 116 O.0287 0.0593 O.3053 1242 0689
Answer #1

HPY of Stock P = (24-18)/18

HPY of Stock P = 0.333

HPY of Stock Q = (116-110)/110

HPY of Stock Q = 0.0545

HPY of Portfolio = 0.75 * 0.333 + 0.25 * 0.0545

HPY of Portfolio = 0.2497 + 0.0136

HPY of Portfolio = 0.2633

Annualized HPY of Portfolio = (1+0.2633)1/3 -1

Annualized HPY of Portfolio = 0.0593

Option B is correct.